Name:

Content and Structure – The paper

1. Has a clearly stated topic or question that it will address – 10

2. Does address the topic or question – 5

3. Has a clearly articulated and logical structure:

a. First paragraph lays out the topic or question – 10 b. Later paragraphs have topic sentences and clearly identifiable main ideas – 10 c. Transitions between paragraphs are logical – 10

4. Accurately presents facts and viewpoints of sources – 10

5. Provides sufficient textual evidence to support the argument – 10

6. Is concise, clear, and not too repetitive – 10

7. Has a clearly identifiable conclusion – 10

8. Has a convincing argument that draws together the evidence presented – 13

Total: 98 (A)

Style and English Usage – The paper

1. Is clear, direct, and appropriate in tone – 20

2. Is free of misspellings – 10

3. Is free of grammatical mistakes – 13

4. Avoids incomplete sentences – 9

5. Avoids run-on or overly complicated sentences – 9

6. Uses correct punctuation – 9

7. Is free of jargon, cliches, and misused words – 9

8. Cites quotations and/or references correctly – 10

Total: 94 (A)

Comments: Muhammad, I still think your very clear and heartfelt reactions to some of the unethical practices described in the book make this a strong essay. As I read it this time, though, it struck me that the idea that an ethical sense is ingrained at birth is at least questionable. There are certainly people who believe that (even from a scientific point of view). However, it seems to me that each of us also forms our sense of ethics through the interactions we have with other people as we grow up. Our parents, teachers, religious leaders, friends, and others all have a part in that. Attitudes that let one person deny the essential humanity of another person are at the root of all of this, and I think that that kind of attitude is always picked up from others.